# Function Encoders: A Principled Approach to Transfer Learning in Hilbert Spaces

### v1

@@ 1.1 Related Work
Transfer learning and generalization have been a key focus of the machine learning community. However, much of the existing literature focuses on statistical shifts (out-of-distribution generalization) rather than structural and task-related perspectives (Zhuang et al., 2021). We employ the definitions from Pan & Yang (2010), which distinguishes transfer learning based on differences between the source and target domains. Our focus is on inductive transfer learning. Several transfer learning approaches focus on modeling task relatedness, including via graph-based representations (Eaton et al., 2008) and manifolds (Ko et al., 2024). We consider a similar geometric view of task relatedness which depends on the relationship between a target tasks and the training data.
@@ 2.1 Inductive Transfer in Hilbert Spaces
Specifically, we focus on an inductive transfer scenario, where the source and target predictors differ, meaning $f_{S}\neq f_{T}$ or, alternatively, $\mathbb{P}(\mathcal{Y}_{S}\mid\mathcal{X}_{S})\neq\mathbb{P}(\mathcal{Y}_{T}\mid\mathcal{X}_{T})$, but the output spaces $\mathcal{Y}_{S}=\mathcal{Y}_{T}$ and domains $\mathcal{D}_{S}=\mathcal{D}_{T}$ are the same. For simplicity of notation, we drop the subscripts ${S}$ and ${T}$ on $\mathcal{X}$ and $\mathcal{Y}$ where appropriate. Additionally, we presume that multiple source datasets $D_{S_{1}},\ldots,D_{S_{n}}$ are available during training, which is similar to the so-called multi-task learning scenario (Caruana, 1997). The model thus has access to multiple datasets during training, and we seek to transfer the knowledge from the source domain $\mathcal{D}_{S}$ and tasks $\mathcal{T}_{S_{1}},\ldots,\mathcal{T}_{S_{n}}$ to a new target task $f_{T}$. This scenario is increasingly common in practice, for instance in image classification, where we have access to a large dataset of diverse images and their corresponding labels and seek to transfer to a new class, or in robotics where we have access to data collected from multiple environments and we seek to transfer to a new environment. Furthermore, we make the following assumption about the tasks.
@@ 2.2 A Geometric View of Transfer
We first consider a typical case of generalization for learned models, where the target predictor lies within the convex hull $H$ of the source predictors $f_{S_{1}},\ldots,f_{S_{n}}$,
$H=\biggl{\{}f\in\mathcal{H}\ \bigg{|}\ f=\sum_{i=1}^{n}\alpha_{i}f_{S_{i}},\sum_{i=1}^{n}\alpha_{i}=1,\alpha_{i}\geq 0\biggr{\}}.$ (1)
@@
Given source predictors $f_{S_{1}},\ldots,f_{S_{n}}$ the target task predictor $f_{T}$ is in the convex hull $H$ of the source predictors.
@@
Given source predictors $f_{S_{1}},\ldots,f_{S_{n}}$ the target task predictor $f_{T}\in\mathrm{span}\{f_{S_{1}},\ldots,f_{S_{n}}\}$. We presume $f\not\in H$ to distinguish extrapolation from interpolation.
@@ 4.1 An Illustrative Polynomial Regression Task
In Figure 2, we see that existing approaches achieve moderate type 1 transfer, interpolating the convex hull. However, existing approaches fail to extrapolate to the linear span (type 2) or to other functions in $\mathcal{H}$ (type 3), while the function encoder using least squares achieves low $L^{2}$ error.
@@ 5 Conclusion
We have introduced a novel categorization of transfer learning based on the geometric interpretation of Hilbert spaces. We made novel improvements to the function encoder algorithm, and argued that it is a natural solution to transfer learning in Hilbert spaces. We empirically validated this argument using four transfer-learning datasets.
@@ Appendix A Basis Function Regularization
As with all (unregularized) least-squares approaches, the Gram matrix can occasionally be close to singular which makes the matrix inverse numerically unstable during training. Thus, we can add a small regularization $\lambda$ to the diagonal of the Gram matrix in the style of Ridge regression, which improves the numerical stability of the inverse. Yet, the regularization term causes the eigenvalues to increase, meaning the solution to the regularized least-squares problem will be a slight under-approximation. Since the basis functions are trained to minimize the error of the approximation, and the approximation is an underestimate, the basis functions constantly grow in magnitude. Thus, we additionally impose an overall regularizer on the diagonal of the Gram matrix to be close to $1$ to prevent this. Note that we do not need the basis functions to be unit, we only require that they do not grow infinitely as training progresses.
@@ Appendix D The Approximate Inner Product and the Distribution of Inputs
We may also have the case that the distribution of inputs and the function change simultaneously. This occurs, for example, in reinforcement learning where the distribution of states is heavily dependent on the hidden parameters. In this case, we still cannot estimate $p$, and so we must use the approximation in (32). However, by doing so we now have a separate inner product defined for each function in the training set. Interestingly, if there exists a set of basis functions which spans all of the training functions, these basis functions are still optimal. Regardless of the inner product, these basis functions would still be able to reproduce the training functions, and therefore minimize the loss. However, if this is not the case, for example if the dimensionality of the space is larger than the number of basis functions, this function-specific inner product is going to affect which solutions are locally optimal. Future work should explore what implications this has, and what changes should be made to account for this problem, if any. Nevertheless, the function encoder can still perform well in at least some circumstances, as demonstrated by the MuJoCo Ant experiment.
@@ Appendix F Architecture
In contrast, parallel neural networks allow basis function individual updates. It is therefore possible for the learning procedure to train only a few basis functions, while the rest remain unused, if the dimensionality of the space is low. We validate this intuition empirically. See Figure 9 for a visualization of this problem.
@@ Appendix G Ablations
We vary the number of basis functions and run FE (LS) on all four datasets. We would expect that using more basis functions improves performance, but there would be diminishing returns. Once the number of basis functions exceeds the dimensionality of the space, performance should stop improving. The experimental results agree with this hypothesis. See Figures 13, 13, 13, and 13. The figures show the performance at the end of training, averaged over the last 50 evaluations. 3 seeds are run per quantity of basis functions, and figures show min, median, and max values over the seeds.
Figure 10: An Ablation on the Number of Basis Functions - Polynomial Dataset. This figures shows the final performance of FE (LS) after training for a varying number of basis functions. The figure shows that the performance initially improves as you increase the number of basis functions, and then levels off.
Figure 11: An Ablation on the Number of Basis Functions - CIFAR Dataset. This figures shows the final performance of FE (LS) after training for a varying number of basis functions. The figure shows steady improvement as the number of basis functions grows, until it levels off around 20 basis functions. This indicates only 20 features are necessary for good performance.
Figure 12: An Ablation on the Number of Basis Functions - 7-Scenes Dataset. This figures shows the final performance of FE (LS) after training for a varying number of basis functions. The figure shows steady improvement as the number of basis functions grows, although it begins leveling off around 100 basis functions. This indicates the dimensionality of the space is relatively high.
Figure 13: An Ablation on the Number of Basis Functions - MuJoCo Ant Dataset. This figures shows the final performance of FE (LS) after training for a varying number of basis functions. For type 1 and type 2 transfer, increasing the number of basis functions improves performance until around 50 basis functions, where it levels off. Interestingly, the number of basis functions does not greatly affect type 3 transfer performance, suggesting type 3 transfer in this setting is challenging.
@@
We vary the number of example data points and evaluate the effect on performance. We would expect that using more example data improves performance, but with diminishing returns. The experimental results, shown in Figures 17, 17, 17, and 17, agree with these claims. The results suggest that a minimal amount of data is required for the function encoder to make accurate predictions, and that the minimal amount depends on the dataset. Similar to the above ablation, the figures show 3 seeds evaluated after training, with their performance averaged over the last 50 evaluation steps.

### v2

@@ 1.1 Related Work
Transfer learning and generalization have been a key focus of the machine learning community. However, much of the existing literature focuses on statistical shifts (out-of-distribution generalization) rather than structural and task-related perspectives (Zhuang et al., 2021). We employ the definitions from Pan & Yang (2010), which distinguishes transfer learning based on differences between the source and target domains. Our focus is on inductive transfer learning. Several transfer learning approaches focus on modeling task relatedness, including via graph-based representations (Eaton et al., 2008) and manifolds (Ko et al., 2024). We consider a similar geometric view of task relatedness which depends on the relationship between a target task and the training data.
@@ 2.1 Inductive Transfer in Hilbert Spaces
Specifically, we focus on an inductive transfer scenario, where the source and target predictors differ, meaning $f_{S}\neq f_{T}$ or, alternatively, $\mathbb{P}(\mathcal{Y}_{S}\mid\mathcal{X}_{S})\neq\mathbb{P}(\mathcal{Y}_{T}\mid\mathcal{X}_{T})$, but the output spaces $\mathcal{Y}_{S}=\mathcal{Y}_{T}$ and domains $\mathcal{D}_{S}=\mathcal{D}_{T}$ are the same. For simplicity of notation, we drop the subscripts ${S}$ and ${T}$ on $\mathcal{X}$ and $\mathcal{Y}$ where appropriate. Additionally, we presume that multiple source datasets $D_{S_{1}},\ldots,D_{S_{n}}$ are available during training, which is similar to the so-called multi-task learning scenario (Caruana, 1997). The model thus has access to multiple datasets during training, and we seek to transfer the knowledge from the source domain $\mathcal{D}_{S}$ and tasks $\mathcal{T}_{S_{1}},\ldots,\mathcal{T}_{S_{n}}$ to a new target task $f_{T}$. This scenario is increasingly common in practice. For instance, in few-shot image classification, we have access to a large dataset of diverse images and their corresponding labels, and we seek to transfer to a new class. In robotics, we have access to data collected from multiple environments, and we aim to transfer dynamics estimates or policies to a new environment. To make use of Hilbert space theory, we make the following assumption about the tasks.
@@ 2.2 A Geometric View of Transfer
We first consider a typical case of generalization for learned models, where the target predictor lies within the convex hull $C_{h}$ of the source predictors $f_{S_{1}},\ldots,f_{S_{n}}$,
$C_{h}=\biggl{\{}f\in\mathcal{H}\ \bigg{|}\ f=\sum_{i=1}^{n}\alpha_{i}f_{S_{i}},\sum_{i=1}^{n}\alpha_{i}=1,\alpha_{i}\geq 0\biggr{\}}.$ (1)
@@
Given source predictors $f_{S_{1}},\ldots,f_{S_{n}}$ the target task predictor $f_{T}$ is in the convex hull $C_{h}$ of the source predictors.
@@
Given source predictors $f_{S_{1}},\ldots,f_{S_{n}}$ the target task predictor $f_{T}\in\mathrm{span}\{f_{S_{1}},\ldots,f_{S_{n}}\}$. We presume $f\not\in C_{h}$ to distinguish extrapolation from interpolation.
@@ 4.1 An Illustrative Polynomial Regression Task
In Figure 2, we see that existing approaches achieve moderate type 1 transfer, interpolation in the convex hull. However, existing approaches fail to extrapolate to the linear span (type 2) or to other functions in $\mathcal{H}$ (type 3), while the function encoder using least squares achieves low $L^{2}$ error.
@@ 5 Conclusion
We have introduced a novel categorization of transfer learning based on the geometric interpretation of Hilbert spaces. We made novel improvements to the function encoder algorithm, and argued that it is a natural solution to transfer learning in Hilbert spaces. We empirically validated this argument using four transfer learning datasets.
@@
Acknowledgments
This work is supported by DARPA HR0011-24-9-0431, NSF 2214939, RTX CW2231110, and AFOSR FA9550-19-1-0005. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding organizations.
@@ Appendix A Basis Function Regularization
As with all (unregularized) least-squares approaches, the Gram matrix can occasionally be close to singular which makes the matrix inverse numerically unstable during training. Thus, we add a small regularization $\lambda$ to the diagonal of the Gram matrix in the style of Ridge regression, which improves the numerical stability of the inverse. Yet, the regularization term causes the eigenvalues to increase, meaning the solution to the regularized least-squares problem will be a slight under-approximation. Since the basis functions are trained to minimize the error of the approximation, and the approximation is an underestimate, the basis functions constantly grow in magnitude. Thus, we additionally impose an overall regularizer on the norm of the basis functions to be close to $1$ to prevent this:
$L_{reg}=\sum_{i=1}^{k}(\lVert g_{i}\rVert^{2}_{\mathcal{H}}-1)^{2}$ (9)
Note that we do not need the basis functions to be unit, we only require that they do not grow infinitely in magnitude as training progresses.
@@ Appendix D The Approximate Inner Product and the Distribution of Inputs
We may also have the case that the distribution of inputs and the function change simultaneously. For example, this occurs in reinforcement learning where the distribution of states is heavily dependent on the hidden parameters. In this case, we still cannot estimate $p$, and so we must use the approximation in (33). However, by doing so we now have a separate inner product defined for each function in the training set. This is true even in the extreme case where the distributions are disjoint, i.e., domain transfer.
Interestingly, even in the presence of distribution shifts, if there exists a set of basis functions which spans all of the training functions, these basis functions are still optimal. Regardless of the inner product, these basis functions would still be able to reproduce the training functions, and therefore minimize the loss. However, if this is not the case, for example if the dimensionality of the space is larger than the number of basis functions, this function-specific inner product is going to affect which solutions are locally optimal. Future work should explore what implications Monte Carlo integration has on performance, and what changes should be made to account for it, if any.
@@ Appendix F Architecture
In contrast, parallel neural networks allow individual basis function updates. It is therefore possible for the learning procedure to train only a few basis functions, while the rest remain unused, if the dimensionality of the space is low. We validate this intuition empirically. See Figure 9 for a visualization of this problem.
@@ Appendix G Ablations
We vary the number of basis functions and run FE (LS) on all four datasets. We would expect that using more basis functions improves performance, but there would be diminishing returns. Once the number of basis functions exceeds the dimensionality of the space, performance should stop improving. The experimental results agree with this hypothesis. See Figures 14, 14, 14, and 14. The figures show the performance at the end of training, averaged over the last 50 evaluations. 3 seeds are run per quantity of basis functions, and figures show min, median, and max values over the seeds. We also consider the cost of increasing the number of basis functions with respect to the training time, and find that it is minor. See Figure 10. Therefore, a reasonable method to choose $k$ is to simply overestimate the dimensionality of the space. We find that $100$ basis functions are enough for many problems.
Figure 10: An Ablation on the Number of Basis Functions vs Compute Time. This figures shows the percentage increase in training time relative to one basis function. The figure shows that the increase in cost is less than 20% for up to 100 basis functions.
Figure 11: An Ablation on the Number of Basis Functions - Polynomial Dataset. This figure shows the final performance of FE (LS) after training for a varying number of basis functions. The performance initially improves as the number of basis functions increases, and then levels off.
Figure 12: An Ablation on the Number of Basis Functions - CIFAR Dataset. This figure shows the final performance of FE (LS) after training for a varying number of basis functions. The figure shows steady improvement as the number of basis functions grows, until it levels off around 20 basis functions. This indicates only 20 features are necessary for good performance.
Figure 13: An Ablation on the Number of Basis Functions - 7-Scenes Dataset. This figure shows the final performance of FE (LS) after training for a varying number of basis functions. The figure shows steady improvement as the number of basis functions grows, although it begins leveling off around 100 basis functions. This indicates the dimensionality of the space is relatively high.
Figure 14: An Ablation on the Number of Basis Functions - MuJoCo Ant Dataset. This figure shows the final performance of FE (LS) after training for a varying number of basis functions. For type 1 and type 2 transfer, increasing the number of basis functions improves performance until around 50 basis functions, where it levels off. Interestingly, the number of basis functions does not greatly affect type 3 transfer performance, suggesting type 3 transfer in this setting is challenging.
@@
We vary the number of example data points and evaluate the effect on performance. We would expect that using more example data improves performance, but with diminishing returns. The experimental results, shown in Figures 18, 18, 18, and 18, agree with these claims. The results suggest that a minimal amount of data is required for the function encoder to make accurate predictions, and that the minimal amount depends on the dataset. Similar to the above ablation, the figures show 3 seeds evaluated after training, with their performance averaged over the last 50 evaluation steps.
@@ Appendix I Baselines
Connections between Siamese networks and Hilbert spaces.
Siamese networks performed well on the few-shot classification problem. We believe this is because there are deep connections between Siamese networks and Hilbert space theory. Siamese networks use a contrastive loss to minimize the distance between inputs belonging to the same category and to maximize the distance otherwise. If you consider the network as a set of basis functions, then the mean output from one category can be interpreted as a mean embedding from kernel literature. Consequently, the difference in these mean embeddings is analogous to the maximum mean discrepancy. Therefore, maximizing the distance between individual embeddings from different classes is similar to maximizing the difference between mean embeddings, and thus Siamese networks are potentially learning basis functions which maximally distinguish between the distributions corresponding to each category.